# Newton numbers, vanishing polytopes and algebraic degrees

### v1

@@ 4.2 $h^{*}$- and $\ell^{*}$-polynomial of agglutination
The $\ell$-Newton number is $\nu^{\ell}_{C}(P)=\frac{1}{2}\ell_{\mathcal{C}}^{*}(\mathcal{P}\sharp_{\mathcal{P}_{B}}\mathcal{P};1)$
By Lemma 3.8, $\ell_{\mathcal{C}}(\mathcal{P}\sharp_{\mathcal{P}_{B}}\mathcal{P},C,F;1)$ is even, so the $\ell$-Newton number is an integer. Recall the definition of the usual Newton number.
@@
(cf. [GKZ94, p.351]) For lattice polytope $P\subset C$ the Newton number is the following alternating sum of volumes:
$\nu_{C}(P)=\sum\limits_{F\leq C}(-1)^{\dim(C)-\dim(F)}Vol_{\mathbb{Z}}(P\cap F)$
@@
    The polynomial $\ell_{\mathcal{C}}^{*}(\mathcal{P}\sharp_{\mathcal{P}_{B}}\mathcal{P};t)$ has nonnegative symmetric coefficients centered at $(n+1)/2$.
@@
    The first item follows from Theorems 3.15, 2.30 and Lemma 2.31;
@@
    The second item follows from Theorems 3.15, 2.30 and Lemma 3.8.
@@
Consider a convenient polytope $P$ in $C=\mathbb{R}^{n}_{\geq 0}$ with face poset $\mathcal{P}$ and boundary poset $\mathcal{P}_{B}$ consisting of its faces in $\partial_{C}P$. Suppose that $T$ is the polyhedral subdivision of $P$ consisting of faces in $\mathcal{P}_{B}$ and their convex hulls with the origin. For $F\in\mathcal{P}_{B}$ denote by $F^{0}$ its convex hull with the origin. The global version of Stapledon’s formulas for monodromy from [Stap17, Corollary 6.14.] implies that:
@@
$\ell_{\mathcal{C}}^{*}(T\sharp_{\mathcal{P}_{B}}\sharp T;t)=\sum\limits_{F\in\mathcal{P}_{B}}(\mathcal{P}_{B},\hat{1}_{\mathcal{C}^{-}},F;t)((t+1)\ell^{*}(F;1)+2\ell^{*}(F^{0};t))$
@@ 5.1 Preliminaries: Mixed volume, Cayley sums, BKK-theorem
Let us recall the mixed volume and Bernstein-Kouchnirenko-Khovansky theory. For an introduction to the convex polyhedral geometry see [Ew], for a survey on this connection between polyhedral geometry and algebraic geometry see [EKK].
@@
Mixed volume $MV$ is the function that maps collections of $n$ convex lattice polytopes in $\mathbb{R}^{n}$ to real numbers and satisfies the following conditions:
@@
    $MV(P,\dots,P)=\operatorname{Vol}_{\mathbb{Z}}(P)$ for any polytope $P$.
@@
    $MV(\lambda P_{1}+\lambda^{\prime}P_{1}^{\prime},P_{2},\dots,P_{n})=\lambda\cdot MV(P_{1},P_{2},\dots,P_{n})+\lambda^{\prime}\cdot MV(P_{1}^{\prime},P_{2},\dots,P_{n}),$
@@
[Ber75] Consider Laurent polynomials $f_{1},\dots,f_{n}\in\mathbb{C}[z_{1}^{\pm 1},\dots,z_{n}^{\pm 1}]$ with Newton polytopes $P_{1},\dots,P_{n}$ and generic coefficients. Then the number of solutions $f_{1}=\dots=f_{n}=0$ in the torus $(\mathbb{C}\setminus 0)^{n}$ is equal to the mixed volume $MV(P_{1},\dots P_{n})$.
@@
$mult(f_{1}\cdot{\dots}\cdot f_{n}\cdot\mathbb{T}^{\Sigma})=\operatorname{Vol}_{\mathbb{Z}}(P)$
@@
([Kh78]; see [Es10, Lemma 1.2] for a proof.) The mixed volume of convex polytopes is zero if and only if some $k$ of these polytopes sum up to a polytope of dimension strictly smaller than $k$.
@@
    there exists $k>0$ such that (after translation):
@@
      $k$ of the polytopes are faces of the same $k$-dimensional volume $1$ lattice simplex in a $k$-dimensional rational subspace $\mathbb{R}^{k}\subset\mathbb{R}^{n}$.
@@
      The images of the other $n-k$ polytopes under the projection $\mathbb{R}^{n}\to\mathbb{R}^{n}/\mathbb{R}^{k}$ have the unit mixed volume.
@@
$\chi(\{f_{1}=\dots=f_{k}=0\})=(-1)^{n-k}\sum\limits_{\begin{subarray}{c}a_{0}>0,\dots,a_{k}>0\\ a_{0}+\dots+a_{k}=n\end{subarray}}MV(\underbrace{P_{0},\dots,P_{0}}_{a_{0}},\dots,\underbrace{P_{k},\dots,P_{k}}_{a_{k}})$
@@
Let $k\leq n$ be non-negative integers. Let $e_{1},\dots,e_{k}$ be the standard basis of $\mathbb{Z}^{k}$. For finite sets $\mathbf{P}_{0},\mathbf{P}_{1},\dots,\mathbf{P}_{k}\subset\mathbb{Z}^{n-k}$, the Cayley sum is defined to be
@@
The volume of the Cayley sum of convex polytopes can be calculated via the mixed volume. The original proofs of these formulas use algebraic geometry (more precisely, they consider the Euler characteristic of certain hypersurfaces in complex tori). These results also admit purely combinatorial proofs using the Cayley trick.
@@
$Vol_{\mathbb{Z}}(P_{0}\ast P_{1}\ast\cdots\ast P_{k})=\sum\limits_{\begin{subarray}{c}a_{0}\geq 0,\dots,a_{k}\geq 0\\ a_{0}+\dots+a_{k}=n-k\end{subarray}}MV(\underbrace{P_{0},\dots,P_{0}}_{a_{0}},\dots,\underbrace{P_{k},\dots,P_{k}}_{a_{k}})$
@@
[Es10, Lemma 1.7.] Consider lattice polytopes $P_{0},\dots,P_{k}\subset\mathbb{R}^{n-k}$. Then we have the following formula
$\sum\limits_{\begin{subarray}{c}I\subset\{1,\dots,k\}\\ I=\{i_{1},\dots,i_{|I|}\}\end{subarray}}(-1)^{k-|I|}Vol_{\mathbb{Z}}(P_{i_{1}}\ast\cdots\ast P_{i_{|I|}})=\sum\limits_{\begin{subarray}{c}a_{0}>0,\dots,a_{k}>0\\ a_{0}+\dots+a_{k}=n-k\end{subarray}}MV(\underbrace{P_{0},\dots,P_{0}}_{a_{0}},\dots,\underbrace{P_{k},\dots,P_{k}}_{a_{k}})$
In other words, we can rewrite the BKK-theorem as certain Newton number. Denote by $C$ the cone $\mathbb{R}^{k}_{\geq 0}\oplus\mathbb{R}^{n}$ and consider any point $Q\in\mathbb{Z}^{n}$, then under assumptions of Theorem 5.7 we have:
@@ 5.2.1 Euler obstructions of sets
Let us recall the formula for the number of points of zero-dimensional critical complete intersections from [Es17, Lemma 2.50]. First, we recall the Euler obstructions of toric varieties from [MT11] (for an introduction to the toric varieties see e.g. [Ful]). Euler obstructions were introduced in [Mcp74] in order to construct characteristic classes of singular complex algebraic varieties. We use the notations from [Es10, Section 1.5].
@@
For a face $\mathbf{A}^{\prime}$ of a finite set $\mathbf{A}\subset\mathbb{R}^{k}$, let $M^{\prime}$ and $M\subset\mathbb{R}^{k}$ be the vector spaces, parallel to the affine spans of $\mathbf{A}^{\prime}$ and $\mathbf{A}$ respectively. We denote the projection $\mathbb{R}^{k}\to\mathbb{R}^{k}/M^{\prime}$ by $\pi_{\mathbf{A}^{\prime}}$. Then the lattice volume of the difference of the convex hulls of $\pi_{\mathbf{A}^{\prime}}(\mathbf{A})$ and $\pi_{\mathbf{A}^{\prime}}(\mathbf{A}\setminus\mathbf{A}^{\prime})$ is denoted by $c^{\mathbf{A}^{\prime}}_{\mathbf{A}}\in\mathbb{Z}$. Set $c^{\mathbf{A}}_{\mathbf{A}}$ to $1$ and $c^{\mathbf{A}^{\prime}}_{\mathbf{A}}$ to 00 if $\mathbf{A}^{\prime}$ is not a face of $\mathbf{A}$.
Consider the square matrix $C$ with entries $c^{\mathbf{A}^{\prime\prime}}_{\mathbf{A}^{\prime}}$, where $\mathbf{A}^{\prime\prime}$ and $\mathbf{A}^{\prime}$ run over the set of all faces of $\mathbf{A}$, and define $e^{\mathbf{A}^{\prime\prime}}_{\mathbf{A}^{\prime}}$ as the $(\mathbf{A}^{\prime\prime},\,\mathbf{A}^{\prime})$-entry of the inverse of $C$. Note that $C$ is upper triangular with $1$’s on the diagonal, if we order faces of $\mathbf{A}$ by their dimension; in particular, the determinant of $C$ equals $1$, and its inverse is integer.
@@
See e.g. [Es10, Example 1.24 and Example 1.26] for an illustration of the combinatorial definitions. Euler obstructions of sets have the following geometric meaning.
@@
[MT11] The Euler obstruction of the set $\mathbf{A}\subset\mathbb{Z}^{k}$ at its face $\mathbf{A}^{\prime}$ equals $(-1)^{\dim\mathbf{A}-\dim\mathbf{A}^{\prime}}$ times the Euler obstruction of the toric variety, corresponding to $\mathbf{A}$, at a point of its orbit, corresponding to $\mathbf{A}^{\prime}$.
@@
A set $\mathbf{A}$ is called smooth at its face $\mathbf{A}^{\prime}$ if $\pi_{\mathbf{A}^{\prime}}(Conv(\mathbf{A}))\setminus\pi_{\mathbf{A}^{\prime}}(Conv(\mathbf{A}\setminus\mathbf{A}^{\prime}))$ is a unimodular simplex.
@@
By the definition of the Euler obstruction, its value is $1$ on the smooth stratum. The projective toric variety corresponding to the set $\mathbf{A}$ is smooth at the orbit corresponding to its face $\mathbf{A}^{\prime}$ if and only if the set $\mathbf{A}$ is smooth at $\mathbf{A}^{\prime}$ and $\mathbf{A}^{\prime}\cap\mathbb{Z}^{n}=\text{aff}(\mathbf{A}^{\prime})\cap\mathbb{Z}^{n}$ (see e.g. [GKZ94, Chapter 5, Theorem 3.16.]). However, it is straightforward to verify (e.g., combinatorially) that $e^{\mathbf{A}^{\prime}}_{\mathbf{A}}=(-1)^{\dim\mathbf{A}-\dim\mathbf{A}^{\prime}}$ for any smooth face $\mathbf{A}^{\prime}$ of $\mathbf{A}$.

### v2

@@ 4.2 $\mathbf{h^{*}}$- and $\ell^{*}$-polynomial of agglutination
The $\ell$-Newton number is defined as $\nu^{\ell}_{C}(P)=\frac{1}{2}\ell_{\mathcal{C}}^{*}(\mathcal{P}\sharp_{\mathcal{P}_{B}}\sharp\mathcal{P};1)$.
By Lemma 3.8, $\ell_{\mathcal{C}}(\mathcal{P}\sharp_{\mathcal{P}_{B}}\sharp\mathcal{P},C,F;1)$ is even, so the $\ell$-Newton number is an integer. Recall the definition of the usual Newton number.
@@
(cf. [GKZ94, p.351]) For a lattice polytope $P\subset C$ the Newton number is the following alternating sum of volumes:
$\nu_{C}(P)=\sum\limits_{F\leq C}(-1)^{\dim(C)-\dim(F)}\operatorname{Vol}_{\mathbb{Z}}(P\cap F).$
@@
    The polynomial $\ell_{\mathcal{C}}^{*}(\mathcal{P}\sharp_{\mathcal{P}_{B}}\sharp\mathcal{P};t)$ has nonnegative symmetric coefficients centered at $(n+1)/2$.
@@
    The first item follows from Theorems 3.15, 2.34 and Lemma 2.35;
@@
    The second item follows from Theorems 3.15, 2.34 and Lemma 3.8.
@@
Consider a convenient polytope $P$ in $C=\mathbb{R}^{n}_{\geq 0}$ with face poset $\mathcal{P}$ and boundary poset $\mathcal{P}_{B}$ consisting of its faces in $\partial_{C}P$. Suppose that $T$ is the polyhedral subdivision of $P$ consisting of faces in $\mathcal{P}_{B}$ and their convex hulls with the origin. For $F\in\mathcal{P}_{B}$ denote by $F^{0}$ its convex hull with the origin. The global version of Stapledon’s formulas for monodromy from [Stap17, Corollary 6.14] implies that:
@@
$\ell_{\mathcal{C}}^{*}(T\sharp_{\mathcal{P}_{B}}\sharp T;t)=\sum\limits_{F\in\mathcal{P}_{B}}\ell_{\mathcal{C}^{-}}(\mathcal{P}_{B},\hat{1}_{\mathcal{C}^{-}},F;t)((t+1)\ell^{*}(F;1)+2\ell^{*}(F^{0};t))$
@@ 5.1 Preliminaries: Mixed volume, Cayley sums, BKK-theorem
Let us recall the mixed volume and Bernstein-Kouchnirenko-Khovanskii theory (BKK-theory in short). For an introduction to the convex polyhedral geometry see [Ew], for a survey on this connection between polyhedral geometry and algebraic geometry see [EKK].
@@
The mixed volume $\operatorname{MV}$ is the unique symmetric multilinear function on $n$-tuples of convex lattice polytopes in $\mathbb{R}^{n}$ that satisfies the following conditions:
@@
    $\operatorname{MV}(P,\dots,P)=\operatorname{Vol}_{\mathbb{Z}}(P)$ for any polytope $P$.
@@
    $\operatorname{MV}(\lambda P_{1}+\lambda^{\prime}P_{1}^{\prime},P_{2},\dots,P_{n})=\lambda\cdot\operatorname{MV}(P_{1},P_{2},\dots,P_{n})+\lambda^{\prime}\cdot\operatorname{MV}(P_{1}^{\prime},P_{2},\dots,P_{n}),$
@@
Definition 5.3.
Consider a polytope $P\subset\mathbb{R}^{n}$ and a covector $\xi\in(\mathbb{R}^{n})^{*}$. The support face $P^{\xi}$ is the maximal face of $P$ on which the linear functional $\xi$ attains its maximum. Consider polynomial $f=\sum\limits_{\omega\in\mathbb{Z}^{n}}a_{\omega}z^{\omega}$ with Newton polytope $N\subset\mathbb{R}^{n}$. For a polytope $\Gamma\subset\mathbb{R}^{n}$ denote by $f|_{\Gamma}$ the restriction $\sum\limits_{\omega\in\Gamma}a_{\omega}z^{\omega}$ of $f$ on $\Gamma$. We denote by $f^{\xi}$ the restriction on the support face $f|_{N^{\xi}}$.
@@
[Ber75] Consider Laurent polynomials $f_{1},\dots,f_{n}\in\mathbb{C}[z_{1}^{\pm 1},\dots,z_{n}^{\pm 1}]$ with Newton polytopes $P_{1},\dots,P_{n}$ and generic coefficients. Then the number of solutions $f_{1}=\dots=f_{n}=0$ in the torus $(\mathbb{C}\setminus 0)^{n}$ is equal to the mixed volume $\operatorname{MV}(P_{1},\dots P_{n})$.
The non-degeneracy conditions are as follows: for any covector $\xi\in(\mathbb{R}^{n})^{*}\setminus 0$ the system $f_{1}^{\xi}={\dots}=f_{n}^{\xi}=0$ has no solutions in the torus $(\mathbb{C}\setminus 0)^{n}$.
@@
$\text{mult}(f_{1}\cdot{\dots}\cdot f_{n}\cdot\mathbb{T}^{\Sigma})=\operatorname{Vol}_{\mathbb{Z}}(P)$
@@
([Kh78]; see [Es10, Lemma 1.2] for a proof). The mixed volume of convex polytopes is zero if and only if some $k$ of these polytopes sum up to a polytope of dimension strictly smaller than $k$.
@@
    there exists $k>0$ such that, after a suitable translation:
@@
      $k$ of the polytopes are faces of the same $k$-dimensional (normalized) volume 11 lattice simplex in a $k$-dimensional rational subspace $\mathbb{R}^{k}\subset\mathbb{R}^{n}$;
@@
      the images of the other nkn-k polytopes under the projection $\mathbb{R}^{n}\to\mathbb{R}^{n}/\mathbb{R}^{k}$ have mixed volume $1$.
@@
$\chi(\{f_{1}=\dots=f_{k}=0\})=(-1)^{n-k}\sum\limits_{\begin{subarray}{c}a_{0}>0,\dots,a_{k}>0\\ a_{0}+\dots+a_{k}=n\end{subarray}}\operatorname{MV}(\underbrace{P_{0},\dots,P_{0}}_{a_{0}},\dots,\underbrace{P_{k},\dots,P_{k}}_{a_{k}})$
@@
Let $k\leq n$ be nonnegative integers. Let $e_{1},\dots,e_{k}$ be the standard basis of $\mathbb{Z}^{k}$. For finite sets $\mathbf{P}_{0},\mathbf{P}_{1},\dots,\mathbf{P}_{k}\subset\mathbb{Z}^{n-k}$, the Cayley sum is defined to be
@@
The volume of the Cayley sum of convex polytopes can be expressed in terms of mixed volumes. The original proofs of these formulas use algebraic geometry (more precisely, they consider the Euler characteristic of certain hypersurfaces in complex tori). These results also admit purely combinatorial proofs using the combinatorial version of the Cayley trick [Stur94, §5].
@@
$\operatorname{Vol}_{\mathbb{Z}}(P_{0}\ast P_{1}\ast\cdots\ast P_{k})=\sum\limits_{\begin{subarray}{c}a_{0}\geq 0,\dots,a_{k}\geq 0\\ a_{0}+\dots+a_{k}=n-k\end{subarray}}\operatorname{MV}(\underbrace{P_{0},\dots,P_{0}}_{a_{0}},\dots,\underbrace{P_{k},\dots,P_{k}}_{a_{k}})$
@@
[Es10, Lemma 1.7] Consider lattice polytopes $P_{0},\dots,P_{k}\subset\mathbb{R}^{n-k}$. Then we have the following formula
$\sum\limits_{\begin{subarray}{c}I\subset\{1,\dots,k\}\\ I=\{i_{1},\dots,i_{|I|}\}\end{subarray}}(-1)^{k-|I|}\operatorname{Vol}_{\mathbb{Z}}(P_{i_{1}}\ast\cdots\ast P_{i_{|I|}})=\sum\limits_{\begin{subarray}{c}a_{0}>0,\dots,a_{k}>0\\ a_{0}+\dots+a_{k}=n-k\end{subarray}}\operatorname{MV}(\underbrace{P_{0},\dots,P_{0}}_{a_{0}},\dots,\underbrace{P_{k},\dots,P_{k}}_{a_{k}})$
In other words, the formula in the BKK-theorem can be expressed as a Newton number. Denote by $C$ the cone $\mathbb{R}^{k}_{\geq 0}\oplus\mathbb{R}^{n-k}$ and consider any point $Q\in\mathbb{Z}^{n}$, then under assumptions of Theorem 5.8 we have:
@@ 5.2.1 Euler obstructions of sets
Let us recall the formula for the number of points of zero-dimensional critical complete intersections from [Es18, Lemma 2.50]. First, we recall the Euler obstructions of toric varieties from [MT11] (for an introduction to the toric varieties see e.g. [Ful]). Euler obstructions were introduced in [Mcp74] for constructing characteristic classes of singular complex algebraic varieties. We use the notations from [Es10, Section 1.5].
@@
For a face $\mathbf{A}^{\prime}$ of a finite set $\mathbf{A}\subset\mathbb{R}^{k}$, let $M^{\prime}$ and $M\subset\mathbb{R}^{k}$ be the vector spaces, parallel to the affine spans of $\mathbf{A}^{\prime}$ and $\mathbf{A}$ respectively. We denote the projection $\mathbb{R}^{k}\to\mathbb{R}^{k}/M^{\prime}$ by $\pi_{\mathbf{A}^{\prime}}$. Then the lattice volume of the set difference $\operatorname{Conv}(\pi_{\mathbf{A}^{\prime}}(\mathbf{A}))\setminus\operatorname{Conv}(\pi_{\mathbf{A}^{\prime}}(\mathbf{A}\setminus\mathbf{A}^{\prime}))$ is denoted by $c^{\mathbf{A}^{\prime}}_{\mathbf{A}}\in\mathbb{Z}$. Set $c^{\mathbf{A}}_{\mathbf{A}}=1$, and set $c^{\mathbf{A}^{\prime}}_{\mathbf{A}}=0$ if $\mathbf{A}^{\prime}$ is not a face of $\mathbf{A}$. These coefficients arise, for instance, in the Gelfand-Kapranov-Zelevinsky decomposition formula (see [GKZ94, Theorem 1.2 from Chapter 10] and [Es10, Proposition 2.10]).
Consider the square matrix $C$ with entries $c^{\mathbf{A}^{\prime\prime}}_{\mathbf{A}^{\prime}}$, where $\mathbf{A}^{\prime\prime}$ and $\mathbf{A}^{\prime}$ run over of all faces of $\mathbf{A}$. Define $e^{\mathbf{A}^{\prime\prime}}_{\mathbf{A}^{\prime}}$ as the $(\mathbf{A}^{\prime\prime},\mathbf{A}^{\prime})$-entry of the inverse matrix $C^{-1}$. Note that $C$ is upper triangular with $1$’s on the diagonal, if we order faces of $\mathbf{A}$ by their dimension; in particular, $\det(C)=1$, and $C^{-1}$ has integer entries.
@@
See e.g. [Es10, Examples 1.24 and 1.26] for illustrations of these combinatorial definitions. Euler obstructions of sets have the following geometric meaning.
@@
[MT11] The Euler obstruction of the set $\mathbf{A}\subset\mathbb{Z}^{k}$ at its face $\mathbf{A}^{\prime}$ equals $(-1)^{\dim\mathbf{A}-\dim\mathbf{A}^{\prime}}$ times the Euler obstruction of the toric variety corresponding to $\mathbf{A}$ at a point in the orbit corresponding to $\mathbf{A}^{\prime}$.
@@
A set $\mathbf{A}$ is called smooth at its face $\mathbf{A}^{\prime}$ if $\pi_{\mathbf{A}^{\prime}}(\operatorname{Conv}(\mathbf{A}))\setminus\pi_{\mathbf{A}^{\prime}}(\operatorname{Conv}(\mathbf{A}\setminus\mathbf{A}^{\prime}))$ is a unimodular simplex.
@@
By the definition of the Euler obstruction, its value is $1$ on the smooth stratum. The projective toric variety corresponding to the set $\mathbf{A}$ is smooth at the orbit corresponding to its face $\mathbf{A}^{\prime}$ if and only if the set $\mathbf{A}$ is smooth at $\mathbf{A}^{\prime}$ and $\mathbf{A}^{\prime}\cap\mathbb{Z}^{n}=\text{aff}(\mathbf{A}^{\prime})\cap\mathbb{Z}^{n}$ (see e.g. [GKZ94, Theorem 3.16 from Chapter 5]). However, one can verify combinatorially that $e^{\mathbf{A}^{\prime}}_{\mathbf{A}}=(-1)^{\dim\mathbf{A}-\dim\mathbf{A}^{\prime}}$ for any smooth face $\mathbf{A}^{\prime}$ of $\mathbf{A}$.